OpenStudy (anonymous):

Use the computational properties of logarithms to write the expression ln (x – 1)(x + 6) in another form.

hero (hero):

Hint: \[\ln(ab) = \ln(a) + \ln(b)\]

OpenStudy (anonymous):

no i was trying to figure the other one out for the other kid